OpenStudy (anonymous):

medal and fan!! Solve for x: −3|2x + 6| = −12

OpenStudy (texaschic101):

-3 | 2x + 6 | = -12 --- divide both sides by -3 | 2x + 6 | = 4 2x + 6 = 4 2x = 4 - 6 2x = -2 x = -1 -(2x + 6) = 4 -2x - 6 = 4 -2x = 4 + 6 -2x = 10 x = -5 x = -5 and x = -1

OpenStudy (anonymous):

thanks ^_^

OpenStudy (texaschic101):

sure thing :)